What is the wavelength in water of a compressional wave whose frequency is 400 per second if the speed of wave in water is 1,450 m/s?


1
Expert's answer
2021-05-10T02:42:04-0400

velocity of wave = frequency "\times" wavelength

"\nu=\digamma\lambda"

1450 m/s = "400 Hz \times \lambda"


wavelenght "\lambda=" "\frac{1450}{400} = 3.625 m"